OpenStudy (pottersheep):

I really need help pleaseee. " Find the dimensions of a rectangle that has a perimeter of 16cm and area of 13 cm^2 "

OpenStudy (pottersheep):

anyone ??:(

OpenStudy (anonymous):

The perimeter of a rectangle can be obtained by the equation \[P = 2(L + W)\] and the area by \[A = LW\] where L and W are different dimensions of the rectangle Do do problems like these, you use one equation to find one dimension in terms of the other, and then substitute that expression into the other equation to find the value of the dimension (and work to find the other dimension as well). I will use the area equation to find the length in terms of width\[L=\frac{13}{W}\] I can substitute that into the equation for the perimeter and then solve for W\[8 = W + \frac{13}{W}\]\[W^2-8W+13=0\]

OpenStudy (anonymous):

Using the quadratic formula, I find the positive root (I do not consider the negative root since the dimensions of the rectangle must be positive). I found\[W = \sqrt{3} + 4\] I know that\[L = \frac{13}{W}\]so I just subtitute the value of W in that equation to find \[L = \frac{13}{\sqrt{3}+4}\]

OpenStudy (pottersheep):

I LOVE YOU !! I GET IT NOW THANKS SOOO MUCH